OpenStudy (anonymous):

A 20.0 kg mass moving at a velocity of + 3.0 m/s is stopped by a constant force of 15.0 N. How many seconds must the force act on the mass to stop it?

OpenStudy (anonymous):

@shamim

OpenStudy (shamim):

Use F=ma 15=20a a=?

OpenStudy (shamim):

Now use v=u-at 0=3-at t=?

OpenStudy (anonymous):

a=.75

OpenStudy (anonymous):

3-2.25= .75 for t as well... or did i do something wrong?

OpenStudy (shamim):

Ok. C v=u-at 0=3-0.75t 0.75t=3 t=3/0.75 t=?

OpenStudy (anonymous):

t=4 thanks again!

OpenStudy (shamim):

U r most welcome!!!!!!!!!!